Fermi National Laboratory
 

Accelerator Update for May 14, 2007 - May 16, 2007

 

Monday May 14

The day shift began with Operations monitoring the Recycler stash (304.8E10), with the Antiproton Source (Pbar) stacking (37.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE experiments taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations prepared CDF for an access at 8:15 AM.

Operations established beam to the Meson MTest experiment at 11:48 AM.

The Neutron Therapy Facility (NTF) reported at 11:53 AM that they had completed treating their last patient for the day.

At 1 PM, Operations reported that the Pbar 95 LCW system supply and return pressures were spiking. Operators investigated and found that an Accumulator quadrupole power supply (A:LQ) was causing the problem. Operations stashed the antiprotons from Pbar to the Recycler and then turned off the power supply; an expert reported that he found a LCW leak in the power supply.

CDF came out of access around 1:30 PM.

Operations put the Tevatron (TeV) through a dry squeeze at 1:55 PM.

Operations put the TeV into shot setup at 3 PM.

Operations reported at 3:59 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations established store 5429 at 5:48 PM with an initial luminosity of 187.13E30.

At 3:48 PM, the Recycler Pelletron tripped off due to a critical device controller (CDC) trip, on a high temperature indication. It reset.

Operations stopped sending beam to the MTest experiment at 6:23 PM.

At 6:28 PM, a Linac Klystron RF station (KRF3) tripped off. Operators investigated and found its breaker tripped. The supply sparked when the operators reset the breaker. Operations called in experts.

A Linac expert repaired KRF3. Linac resumed sending beam at 10:09 PM.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Tuesday May 15

The midnight shift began with Operations monitoring store 5429 (93.21E30) and the stash (133.1E10), with Pbar stacking (27.9E10), and with NuMI and MiniBooNE taking beam.

At 2:53 AM, a Booster RF station (BRF6) tripped off for the first of three times in about three hours. It reset.

At 3:39 AM, the Pelletron’s VME crate wasn’t responding to a node poll. Operators power cycled the crate, but it did not help. Operations contacted an expert.

At 5:54 AM, BRF6 tripped off. It reset.

Operations established beam to MTest at 6:15 AM.

At 7:41 AM, the Pelletron suffered a full discharge. The expert was here and took care of it.

A Recycler expert brought the Pelletron back on line at 7:58 AM, but then it tripped off five minutes later. It reset.

Operations reported at 8:09 AM that the Recycler had stashed twice during the midnight shift.

Operations terminated store 5429 at 9:04 AM.

Operations prepared CDF and D0 for access at 9:15 AM.

Cryo system technicians began repairing a TeV sector A1 wet engine at 10:15 AM.

A MI expert began a study at 10:29 AM.

At 11:08 AM, the Pelletron suffered another full discharge. Experts investigated.

DO came out of access at 1:25 PM

Recycler experts brought the Pelletron back online at 1:37 PM.

CDF came out of access at 1:49 PM.

Operations put the TeV through a dry squeeze at 2:09 PM.

Operations put the TeV through a wet squeeze at 3:01 PM.

MTest reported at 3:05 PM that they had completed their test run and would be taking the experiment apart. Operations disabled the beam line.

At 3:15 PM, the I- Source tripped off. Experts investigated and found that the Haefely anode power supplies had been turned down so that a quadrupole (Q3) would run better, but this reduced the temperature inside the supply. Due to the roll up doors being open and the humidity, water condensed on the copper tubing inside the power supplies and caused an arc. Techs dried and cleaned the carbon off of the supply. Beam resumed at 3:57 PM.

Operations put the TeV into shot setup at 3:57 PM.

Operations reported at 4:12 PM that the Recycler had stashed once during the day shift.

At 4:23 PM, Linac experts reported that the H- Source, which had tripped off when the I- Source had tripped, was out of commission. This will be further investigated on Wednesday.

Operations established store 5432 at 6:15 PM with an initial luminosity of 202.77E30.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Wednesday May 16

The midnight shift began with Operations monitoring store 5432 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:30 AM, BRF6 tripped off for the first of two times in three and a half hours. It reset.

Operations terminated store 5432 at 5:59 AM.

Pbar experts began reverse proton studies at 6:23 AM.

Operations prepared CDF for access at 6:23 AM.

Operations began preparing the TeV for access at 6:44 AM.

Recycler experts began conditioning the Pelletron at 7:56 AM.

Operations reported at 8 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to allow CDF to investigate and hopefully make repairs to their silicon detector cooling system. The investigation and repairs might take 16 hours or more. The Run Coordinator will allow accelerator maintenance and studies up to 4 PM today.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/16/2007    email Fermilab